OpenStudy (luigi0210):

Determine:

OpenStudy (luigi0210):

\[\int\limits_{0}^{1} \frac{x+1}{x^2+2x} dx\]

zepdrix (zepdrix):

\[\large u=x^2+2x \qquad\to\qquad du=(2x+2)dx\qquad\to\qquad du=2(x+1)dx\] \[\Large \int\limits_0^1\frac{(x+1)dx}{x^2+2x}\]

zepdrix (zepdrix):

Woops I still needed to divide by 2 to match up our (x+1)dx. Understand how that works? :o

OpenStudy (luigi0210):

So far, yes.

OpenStudy (luigi0210):

If I did it right, the answer should be DNE

zepdrix (zepdrix):

From that last step, dividing each side by 2 gives us,\[\large \color{purple}{\frac{1}{2}du=(x+1)dx}\]\[\Large \int\limits\limits_0^1\frac{\color{purple}{(x+1)dx}}{x^2+2x} \qquad\to\qquad \int\limits\limits_{x=0}^1\frac{\color{purple}{\dfrac{1}{2}du}}{u}\] DNE? Hmm let's see..

zepdrix (zepdrix):

\[\large \frac{1}{2}\ln|x^2+2x|\quad|_0^1\] Oh cause of that pesky 0? :3 I see..

OpenStudy (luigi0210):

Yes, I just wanted to make sure I was right. thanks :)

zepdrix (zepdrix):

Yah it appears this does not converge :O Good job!

OpenStudy (luigi0210):

@Loser66 has something to say

OpenStudy (loser66):

to me, I break the problem into 2 parts: